California's coastline is 106 miles longer than twice the coastline of Texas. It is also 473 miles longer than the coastline of Texas. Find the lengths of the coastlines of California and Texas.
Answer by Gogonati(855)

Coastline of Texas: x
Coastline of California's: 2x+106
Coastline of Cal. x+473
Solve the equation: x+473=2x+106 x=472-106 x=367.
Cal. coastline: 367+473=840
Tex. coastline:367